# High-throughput Parasitic-independent Probe Thermal Resistance Calibration for Robust Thermal Mapping with Scanning Thermal Microscopy

### v2

@@ Abstract
Nanostructured materials, critical for thermal management in semiconductor devices, exhibit a strong size dependence in thermal transport. Specifically, studying the variation of thermal resistance across grain boundaries is of critical importance in designing effective thermal interface materials for heterogeneous integration. Frequency-domain Thermoreflectance (FDTR)-based techniques can provide thermal resistance mapping in the micrometer (m) length scale. Scanning Thermal Microscopy (SThM) has the potential for quantification of local thermal transport with orders of magnitude higher spatial resolution (<100 nm). However, challenges in the accurate quantification of the raw signal to thermal conductivity and surface sensitivity limit its widespread adoption as a characterization standard for understanding nanoscale heat transport and defect mediated thermal properties modulation in nanostructured films. Here, we introduce a circuit-based probe thermal resistance () calibration technique that is independent of any parasitic heat transport pathways; hence, using the calibration allows us to accurately measure the heat dissipated from the SThM probe and the resulting tip temperature change to extract the thermal resistance of the film (). Following calibration, SThM achieved sub-100 nm spatial resolution in mapping thermal resistance across a 15 nm-thick Al film deposited via e-beam evaporation on SiO2 substrate. Finally, the thermal resistance mapping is converted to thermal conductivity using the finite element modeling based calibration technique, where the average of the pixel-level values yields the effective thermal conductivity of 2.87 0.18 W m-1K-1, in good agreement with published theoretical frameworks explaining heat transport modes in ultra-thin Al.
@@ I Introduction
Across the last two decades, continuous materials scaling coupled with precise nanostructuring has powered a remarkable run of progress in electronic, photonic, and quantum technologies, reinforced by advances in MEMS/NEMS, nanoscale sensors for biomedicine, and scanning-probe/optical instrumentation. Yet the very act of miniaturization reshapes transport, mechanics, and interfacial phenomena, making size-dependent property shifts an essential focus of modeling, metrology, and design. For example, silicon (Si) thermal conductivity () is reduced from its bulk value of 138 to 2.2 for a 155 nm Si thin filmmcconnell2005thermal. When the physical dimension becomes comparable to or smaller than the mean free path () of heat carriers (electrons and phonons), the macroscopic equilibrium and continuum assumptions break down. Therefore, understanding the nanoscale energy transport phenomena, both analytically and experimentally, is crucial for the continued miniaturization of technology.
@@
Thermal metrology plays a vital role in validating theoretical models and advancing the understanding of energy transport mechanisms. Conventional techniques such as -thermocouples, microbridge and suspended devices (), infrared thermography (IR), and liquid crystal thermography are used to measure macroscopic thermal properties of devices. Recently, with technological advancements in device fabrication and deposition processes, the technique has been extended for use with thin films and microscale devices by fabricating bridges a few thick cahill1987thermal. This method enables accurate characterization of both in-plane and cross-plane thermal transport in metallic, semiconducting, and dielectric thin films. A major limitation of the 3 technique is its requirement for complex sample preparation and the complexities arising from voltage oscillations when measuring metals and semiconductors, thus limiting its applicability to a narrow range of materialscahill2002thermometry. Optical based techniques such as frequency- and time-domain thermoreflectance spectroscopy (FDTR and TDTR), Raman spectroscopy, laser flash techniques, and fluorescence techniques are widely used for measuring the temperature and thermal transport properties of microstructured materials christofferson2008microscale. In contrast to the method, optical-based techniques such as TDTR and FDTR enable non-contact characterization of thermal properties, minimizing physical interference with the sample. Recent advancements in laser source technology have significantly enhanced the capability of TDTR and FDTR techniques to characterize the thermal properties of materials with superior temporal and spatial resolution. Depending on the type of laser used in the measurement, these methods can achieve temporal resolutions ranging from a few nanoseconds to 30 fs (1 fs = sec). It is worth noting that the electronic relaxation time () typically lies in the range of about 10–50 fs. Therefore, TDTR and FDTR are useful techniques for measuring the thermal transport properties associated with electron-phonon coupling mechanisms. Conventional TDTR is laterally resolution-limited by the pump/probe spot size—typically 0.5–10 and therefore cannot reliably image thermal variations arising from nanoscale defects.
@@
In the late 1990s, the development of scanning probe microscopy (SPM)–based techniques, such as scanning thermal microscopy (SThM) and near-field scanning optical microscopy (NSOM) provided unprecedented spatial resolution for nanoscale characterization and manipulation. In general, NSOM requires optically compatible and flat samples, is limited by weak signal-to-noise ratios, and exhibits greater tip fragility compared to SThM. Current commercially available SThMs are based on the working principle designed by Majumdar et al.majumdar1993thermal as illustrated in Fig. 1(a). The SThM system uses specialized thermal probes (TPs) with tip radii below 100 nm, capable of resolving temperature differences smaller than 0.1 with sub-100 nm spatial resolution, thereby enabling quantitative nanoscale thermal characterization of materials. SThM can operate in either temperature contrast mode (TCM) or conductivity contrast mode (CCM). In TCM, a tip-integrated thermocouple measures the junction temperature at the tip-sample interface, whereas the CCM mode provides qualitative mapping of local thermal conductivity across a broad range of materials, including metallic, semiconducting, and dielectric thin films. Transforming these qualitative signals into quantitative thermal conductivity values is not trivial because of the complex coupled thermal transport at the tip-sample interface. There is no standardized technique for obtaining quantitative measurements of thermal resistance with SThM. Recently, various theoretical modeling, calibration strategies, and instrument add-on techniques have been reportedgucmann2021scanning; li2022comprehensive; li2025quantitative. Bodzenta et al. and Li et al. provided comprehensive reviews of SThM theoretical models and experimental strategies for quantitative thermal measurementsbodzenta2022scanning; li2025quantitative.
@@
In this work, we present a circuit-based calibration technique for measuring probe thermal resistance (). Compared with conventional thermal-stage methods, this calibration approach reduces sensitivity to temperature-controller errors and environmental factors (e.g., moisture), and is less affected by common artifacts, including calibration-sample variations. First, we construct thermal-resistance maps across the film surface by coupling scanning thermal measurements with finite-element heat-transfer modeling of the tip–sample mechanical contact. These maps resolve spatial variations in thermal transport that arise from differences in crystallite size and orientation, effects that are most pronounced in polycrystalline and amorphous films. Because the maps carry the imprint of surface preparation and condition—the same steps a film undergoes in routine device-integration flows—the resulting resistance estimates reflect the as-processed device film. Building on the same calibrated contact model, we then extract the films’ effective thermal conductivity, , using an experimentally anchored procedure that combines accurate instrument calibration with contact heat-transfer modeling. Section II of this paper discusses the theoretical modeling of SThM required for understanding different heat transfer mechanisms at the tip-sample contact. Section III describes the methodology used to extract of samples and the post-processing steps to study the underlying nanoscale thermal transport physics and to map the corresponding thermal resistances. In Section IV, we present the results obtained through the techniques described in Sections II and III. Finally, we summarize our results with concluding remarks.
@@ II Background of The Quantitative SThM Measurement Technique
Theoretical modeling links the electrical signals obtained from a lock-in amplifier or DC detector to the thermal properties of the sample. Accurate quantitative measurements of thermal conductivity heavily rely on mathematical modeling to transform probe voltage signals into thermal conductivity values. Thermal transport between the probe and sample involves multiple heat transfer channels, including conduction through the solid–solid contact (tip–sample) and the water meniscus (from the moisture), convection through air, and thermal radiation, which becomes significant at nanometer scale tip–sample distances (see Fig. 2). Each of these channels has its corresponding contribution, which depends on the instrumentation, measurement environment, and sample under consideration. Various analytical and numerical approaches have been reported for studying each heat transfer channel and its significance in quantitative measurements. A recent review by Li et al. comprehensively discusses multiple analytical and numerical approaches for calculating heat transfer through these channels li2025quantitative.
@@
This work uses a thermistor probe tip that consists of Palladium (Pd) resistor material with a thermal coefficient of resistance (TCR) of 3800 . The sensor temperature is measured by applying an electric current to the probe, and the drop in voltage across the probe is measured with the help of the Wheatstone bridge, as shown in Fig.1(a). Since Pd is a metal, within a relatively small temperature range, the electrical resistance of the tip (R) varies linearly with temperature (T).
where is the resistance of the tip at temperature and is the TCR.
@@
When the probe is in air without contacting the sample, heat is dissipated by conduction through the cantilever, convection through the air, and radiation to the ambient, as shown in Fig. 2(a). The heat transport is governed by Fourier’s law of conduction, as presented in Eq. (2). When the tip contacts the sample, solid tip–sample contact adds a heat-conduction path, represented by Eq. (3). The steep tip–sample temperature gradient may challenge the applicability of Fourier’s law. However, if the dominant heat carriers have a mean free path () much smaller than the tip radius (50 nm), and the slow scan rate (>1 ms per pixel) maintains local thermal equilibrium, Fourier’s law with an effective thermal conductivity can still be applied as a simple analytical model of heat transport. More discussion on this will be found in the methods section. Based on the heat-transfer channels illustrated in Fig. 2, the relationship between temperature and thermal resistance is expressed as follows:
@@
The current applied to the probe through the integrated SThM power controller induces Joule heating in both the probe and the tip material. The probe current in CCM is higher than that in TCM. Under the constant-current condition of CCM, the generated electrical heat () flows from the tip into the sample. Therefore, the amount of heat transfer depends on the thermal conductivity of the sample. For highly conductive materials, heat spreads quickly, leading to a small temperature rise in the sensor and a low output voltage. In contrast, for materials with high thermal resistance, heat accumulates near the tip, resulting in a larger temperature rise and a higher voltage output. The corresponding temperature change () can be expressed as shown in Eq. (6). Experimental verification of this relation is provided in Fig. S3 (Supplementary Material).
@@
Furthermore, since the tip radius (b) is very small, the associated thermal healing radius () is negligible relative to the sample size (>). Thus, under thermal equilibrium gu2002imaging, it is reasonable to approximate . Consequently, we obtain the following expression:
Therefore, using Eq. (10), the values of and can be expressed in terms of the measured thermal signals ( and ). Here, embodies the sample’s thermal transport characteristics, and its detailed formulation will be discussed in the Methods section.
@@ III.2 Experimental setup and numerical data generation
The thermal signal data are obtained from the experiments conducted using an AFM (Dimension Icon® manufactured by Bruker Inc.) equipped with an integrated SThM module. The SThM probe is purchased from Bruker Corp. and it consists of a thin Pd line on Si3N4 cantilever, as shown in Fig. S5 (Supplementary Material). This probe is pre-mounted on the probe holder, and the probe tip serves as one of the arms of the Wheatstone bridge, as shown in the schematic in Fig. 1(a). Before each SThM scan, the Wheatstone bridge is balanced to minimize background thermal signal noise. Subsequently, an electrical voltage (> 0.5 V) is applied to the probe. The temperature increase at the SThM tip modifies the resistance of the Pd heater, producing a voltage imbalance across the Wheatstone bridge. This imbalance is detected as the SThM voltage signal (). As discussed in Sec. II, an increase in reflects a higher SThM probe resistance and, correspondingly, a higher tip temperature, which arises under conditions of low local thermal conductance of the sample. All measurements were conducted at a constant contact setpoint force and scan rate to ensure consistency. We used a gain of while extracting the change in voltage of the Pd heater. This gain helps amplify the corresponding thermal signals without any loss of information.
For thermal relaxation and steady-state measurements, the SThM tip rasters at a relatively low scan rate of 0.7 Hz. This scan rate is maintained throughout the experiment to reduce variance in thermal signals due to thermal drift. The acquired images are analyzed using Gwyddion software, while the pixel-wise thermal signal data are extracted and processed through a specially developed Python code tailored for this study.
@@ III.3 Measuring probe thermal resistance ()
We propose a circuit-based technique to measure the non-contact thermal resistance () by monitoring the probe-tip Joule heating power and the resulting temperature rise under an applied voltage. Unlike stage-based techniques, this approach eliminates dependence on thermal stage sensitivity and minimizes the influence of parasitic thermal effects from the environment. In this approach, Joule heating from the applied current raises the tip temperature. The temperature rise is directly proportional to the input power, where the proportionality constant corresponds to the non-contact thermal resistance, (Eq.11). The temperature rise () was calculated relative to the initial tip temperature, which was at room temperature ( = 294.2 K), and the corresponding initial electrical resistance of the tip ( = 177.2 ). It is important to note that the probe electrical resistance (), which is measured using a Wheatstone bridge, differs from the tip resistance () due to the presence of built-in limiting resistors in the probe, as shown in Fig. S4 (Supplementary Material). The applied power () was determined directly from Joule heating in the probe under the applied current. The resulting data were fitted using linear regression, yielding an value of K/W as shown in Fig. 3. An instrumental uncertainty of K in was considered during the analysis. This result is consistent with previously reported values for similar probes, thereby validating the accuracy of the proposed measurement technique huang2024violation; razeghi2023single.

### v3

@@ Abstract
Nanostructured materials, critical for thermal management in semiconductor devices, exhibit a strong size dependence in thermal transport. Specifically, studying the variation of thermal resistance across grain boundaries is of critical importance in designing effective thermal interface materials for heterogeneous integration. Frequency-domain Thermoreflectance (FDTR)-based techniques can provide thermal resistance mapping in the micrometer (m) length scale. Scanning Thermal Microscopy (SThM) has the potential for quantification of local thermal transport with orders of magnitude higher spatial resolution (< 100 nm). However, challenges in the accurate quantification of the raw signal to thermal conductivity and surface sensitivity limit its widespread adoption as a characterization standard for understanding nanoscale heat transport and defect-mediated thermal properties modulation in nanostructured films. Here, we introduce a circuit-based probe thermal resistance () calibration technique that is independent of any parasitic heat transport pathways; thus, the calibration allows accurate determination of the heat dissipated from the SThM probe and the resulting tip temperature increase, thereby enabling extraction of the local thermal resistance of the film. In this study, SThM achieved sub-100 nm spatial resolution in mapping thermal resistance across a 15 nm-thick Al film deposited via e-beam evaporation on SiO2 substrate. Finally, the thermal resistance mapping is converted to thermal conductivity using the finite element modeling-based calibration technique, where the average of the pixel-level values yields the effective thermal conductivity of 2.87 0.18 W/mK, in good agreement with published theoretical frameworks explaining heat transport modes in ultra-thin Al.
@@ I Introduction
Across the last two decades, continuous materials scaling coupled with precise nanostructuring has powered a remarkable run of progress in electronic10019538; munde20253d; banerjee20023, photonicostroverkhova2016organic; huang2020recent, and quantum technologiesputranto2024deep, reinforced by advances in MEMS/NEMS zhu2019development; lyshevski2018mems, nanoscale sensors for biomedicinehannah2020developments; wang2013advances; tovar2023recent, and scanning-probe/optical instrumentationcahill2003nanoscale. Miniaturization fundamentally modifies transport, mechanical, and interfacial behavior. As a result, size-dependent variations in material properties become critical considerations in modeling, metrology, and device design. For example, silicon (Si) thermal conductivity () is reduced from its bulk value of 138 W/mK to 2.2 W/mK for a 155 nm Si thin filmmcconnell2005thermal; huang2019thermal. When the physical dimension becomes comparable to or smaller than the mean free path () of heat carriers (electrons and phonons), the macroscopic equilibrium and continuum assumptions break down. Therefore, understanding the nanoscale energy transport phenomena, both analytically and experimentally, is crucial for the continued miniaturization of technology.
@@
Thermal metrology plays a vital role in validating theoretical models and advancing the understanding of energy transport mechanisms. Conventional techniques such as -thermocouples, microbridge and suspended devices (), infrared thermography (IR), and liquid crystal thermography are used to measure macroscopic thermal properties of devices. Recently, with technological advancements in device fabrication and deposition processes, the technique has been extended for use with thin films and microscale devices by fabricating bridges a few thick cahill1987thermal. This method enables accurate characterization of both in-plane and cross-plane thermal transport in metallic, semiconducting, and dielectric thin films. A major limitation of the 3 technique is its requirement for complex sample preparation and the complexities arising from voltage oscillations when measuring metals and semiconductors, thus limiting its applicability to a narrow range of materialscahill2002thermometry. Optical based techniques such as frequency- and time-domain thermoreflectance spectroscopy (FDTR and TDTR), Raman spectroscopy, laser flash techniques, and fluorescence techniques are widely used for measuring the temperature and thermal transport properties of microstructured materials christofferson2008microscale; tiwari2013anisotropic; zeng2017photothermal. In contrast to the method, optical-based techniques such as TDTR and FDTR enable non-contact characterization of thermal properties, minimizing physical interference with the sample. Recent advancements in laser source technology have significantly enhanced the capability of TDTR and FDTR techniques to characterize the thermal properties of materials with superior temporal and spatial resolution isotta2023microscale. Depending on the type of laser used in the measurement, these methods can achieve temporal resolutions ranging from a few nanoseconds to 30 fs (1 fs = sec). It is worth noting that the electronic relaxation time () typically lies in the range of about 10–50 fs. Therefore, TDTR and FDTR are useful techniques for measuring the thermal transport properties associated with electron-phonon coupling mechanisms. However, these optical-based thermal measurement methods suffer from an intrinsic spatial resolution limitation imposed by the pump–probe spot size (typically 0.5–10 m), rendering them inadequate for reliably imaging thermal heterogeneities arising from nanoscale defects.
@@
In the late 1990s, the development of scanning probe microscopy (SPM)–based techniques, such as scanning thermal microscopy (SThM) and near-field scanning optical microscopy (NSOM) provided unprecedented spatial resolution for nanoscale characterization and manipulation. In general, NSOM requires optically compatible and flat samples, is limited by weak signal-to-noise ratio, and exhibits greater tip fragility compared to SThM. Table 1 summarizes the quantitative capabilities of thermal metrology techniques for microelectronics. Current commercially available SThMs are based on the working principle designed by Majumdar et al.majumdar1993thermal as illustrated in Fig. 1(a). The SThM system uses specialized thermal probes (TPs) with tip radii below 100 nm, capable of resolving temperature differences smaller than 0.1 with sub-100 nm spatial resolution, thereby enabling quantitative nanoscale thermal characterization of materials. SThM can operate in either temperature contrast mode (TCM) or conductivity contrast mode (CCM). In TCM, a tip-integrated thermocouple measures the junction temperature at the tip-sample interface, whereas the CCM mode provides qualitative mapping of local thermal conductivity across a broad range of materials, including metallic, semiconducting, and dielectric thin films. Transforming these qualitative signals into quantitative thermal conductivity values is not trivial because of the complex coupled thermal transport at the tip-sample interface. There is no standardized technique for obtaining quantitative measurements of thermal resistance with SThM. Recently, various theoretical modeling, calibration strategies, and instrument add-on techniques have been reportedgucmann2021scanning; li2022comprehensive; li2025quantitative. Bodzenta et al. and Li et al. provided comprehensive reviews of SThM theoretical models and experimental strategies for quantitative thermal measurementsbodzenta2022scanning; li2025quantitative.
@@
In this work, we present a circuit-based calibration technique for measuring non-contact thermal resistance (), also known as probe thermal resistance. Compared with conventional thermal-stage methods, this calibration approach reduces sensitivity to temperature-controller errors and environmental factors (e.g., moisture), and is less affected by common artifacts, including calibration-sample variations. First, we construct thermal-resistance maps across the film surface by coupling scanning thermal measurements with finite-element heat-transfer modeling of the tip–sample mechanical contact. These maps resolve spatial variations in thermal transport that arise from differences in crystallite size and orientation, effects that are most pronounced in polycrystalline and amorphous films. Because the maps carry the imprint of surface preparation and condition—the same steps a film undergoes in routine device-integration flows—the resulting resistance estimates reflect the as-processed device film. Building on the same calibrated contact model, we then extract the film’s effective thermal conductivity () using an experimentally anchored procedure that combines accurate instrument calibration with contact heat-transfer modeling. Sec. II of this paper discusses the theoretical modeling of SThM required for understanding different heat transfer mechanisms at the tip-sample contact. Sec. III describes the methodology used to extract of samples and the post-processing steps to study the underlying nanoscale thermal transport physics and to map the corresponding thermal resistances. In Sec. IV, we present the results obtained through the techniques described in Sec.s II and III. Finally, we summarize our results with concluding remarks.
| Technique | Mechanism | Spatial Resolution | Temperature Resolution |
| Infrared thermography | Thermal radiation (Planck, Stefan–Boltzmann) | 3–10 m | 0.02–1 K |
| Photothermal Raman | Raman shift or Stokes/anti-Stokes ratio | 0.5–1 m | 1–10 K |
| 3 method | AC Joule heating with electrical sensing | — | 0.1 mK–1 K |
| TDTR / FDTR | Reflectivity-based transient thermometry | 1–10 m | 0.01 K |
| NSOM | Near-field optical heat detection | 50–100 nm | 0.1–1 K |
| SThM | AFM-integrated thermocouple/Pd thermistor tip | 10–50 nm | 0.1 K |
Data adapted from Refs. christofferson2008microscale; cahill1987thermal; zhang2007nano.
@@ II Background of The Quantitative SThM Measurement Technique
Theoretical modeling links the probe voltage signals measured by the Wheatstone bridge detector to the sample’s thermal properties. Therefore, accurate quantitative measurements of thermal conductivity heavily rely on mathematical modeling to transform probe voltage signals into thermal conductivity values. Thermal transport between the probe and sample involves multiple heat transfer channels, including conduction through the solid–solid contact (tip–sample) and the water meniscus (from the moisture), convection through air, and thermal radiation, which becomes significant at nanometer scale tip–sample distances (see Fig. 2). Each of these channels has its corresponding contribution, which depends on the instrumentation, measurement environment, and sample under consideration. Various analytical and numerical approaches have been reported for studying each heat transfer channel and its significance in quantitative measurements. A recent review by Li et al. comprehensively discusses multiple analytical and numerical approaches for calculating heat transfer through these channels li2025quantitative.
@@
This work uses a thermistor probe tip that consists of Palladium (Pd) resistor material with a thermal coefficient of resistance (TCR) of 3800 , as specified by the manufacturer. Importantly, the TCR value is not affected by the experimental procedure or daily use of the probepernot2021frequency; huang2024violation. Since Pd is a metal, within a relatively small temperature range, the electrical resistance of the tip () varies linearly with temperature (T).
where is the electrical resistance of the tip at temperature and is the TCR. The tip temperature is measured by applying an electric current to the probe and monitoring the change in electrical resistance of the tip with SThM software.
@@
When the probe is in air without contacting the sample, heat is dissipated by conduction through the cantilever, convection through the air, and radiation to the ambient, as shown in Fig. 2(a). The heat transport is governed by Fourier’s law of conduction, as presented in Eq. (2). The steep tip–sample temperature gradient may challenge the applicability of Fourier’s law. However, when the mean free path () of the dominant heat carriers is much smaller than the tip radius (50 nm), and the SThM operates at slow scan rates (0.5–1 Hz, corresponding to ms per pixel), local thermal equilibrium is maintained.deshmukh2022direct; swoboda2023spatially. Therefore, Fourier’s law with an effective thermal conductivity can still be applied as a simple analytical model of heat transport. When the tip contacts the sample, a solid–solid tip–sample junction introduces an additional heat-conduction pathway, as described by Eq. (3). Based on the heat-transfer channels illustrated in Fig. 2(a,b), the relationships between the thermal resistances in non-contact and contact configurations are given by Eqs. (4) and (5), respectively.
@@
The current applied to the probe through the integrated SThM power controller induces Joule heating in both the probe and the tip material. The probe current in CCM is higher than that in TCM. Under the constant-current condition of CCM, the generated electrical heat () flows from the tip into the sample. Therefore, the amount of heat transfer depends on the thermal conductivity of the sample. For highly thermally conductive materials, heat dissipates rapidly, resulting in a small temperature rise at the tip and a correspondingly low output voltage. In contrast, materials with high thermal resistance impede heat flow, causing heat to accumulate near the tip and producing a larger temperature rise and higher voltage output. The corresponding temperature change () can be expressed as shown in Eq. (6). Experimental verification of this relation is provided in Fig. S3 (Supplementary Material).
@@
Furthermore, since the tip radius (b) is very small, the associated thermal radius () is negligible relative to the sample size (>). Thus, under thermal equilibrium gu2002imaging, it is reasonable to approximate . Consequently, we obtain the following expression:
Therefore, using Eq. (10), the values of and can be expressed in terms of the measured thermal signals ( and ). Here, embodies the sample’s thermal transport characteristics, and its detailed formulation will be discussed in Sec. III.
@@ III.2 Experimental setup and numerical data generation
The thermal signal data are obtained from the experiments conducted using an AFM (Dimension Icon® manufactured by Bruker Inc.) equipped with an integrated SThM module. The SThM probe is purchased from Bruker Corp. and it consists of a thin Pd line on Si3N4 cantilever, as shown in Fig. S5 (Supplementary Material). This probe is pre-mounted on the probe holder, and the probe tip serves as one of the arms of the Wheatstone bridge, as shown in the schematic in Fig. 1(a). Before each SThM scan, the Wheatstone bridge is balanced to minimize background thermal signal noise. Subsequently, an input voltage (> 0.5 V) is applied to the probe. The temperature increase at the SThM tip changes the resistance of the Pd heater, producing a voltage imbalance across the Wheatstone bridge. This imbalance is detected as the SThM voltage signal (). All measurements were conducted at a constant contact setpoint force and scan rate to ensure consistency. We used a gain of while extracting the change in voltage of the Pd heater. This gain helps amplify the corresponding thermal signals without any loss of information.
For thermal relaxation and steady-state measurements, the SThM tip rasters at a relatively low scan rate of 0.5–1 Hz (S2 supplementary material). This scan rate is maintained throughout the experiment to reduce variance in thermal signals due to thermal drift. The acquired images are analyzed using Gwyddion software, while the pixel-wise thermal signal data are extracted and processed through a specially developed Python code tailored for this study.
@@ III.3 Measuring probe thermal resistance ()
We propose a circuit-based technique to measure the non-contact thermal resistance () by monitoring the probe-tip Joule heating power and the resulting temperature rise under an applied voltage. Unlike stage-based techniques, this approach eliminates dependence on thermal stage sensitivity and minimizes the influence of parasitic thermal effects from the environment. In this approach, Joule heating from the applied current raises the tip temperature. The temperature rise is directly proportional to the input power, where the proportionality constant corresponds to the non-contact thermal resistance, (Eq.12). The temperature rise () was calculated relative to the initial tip temperature, which was at room temperature ( K), and the corresponding initial electrical resistance of the tip ( 177 ). It is important to note that the probe electrical resistance (), which is output of balanced Wheatstone bridge, differs from the tip resistance () due to the presence of built-in limiting resistors in the probe, as shown in Fig. S4 (Supplementary Material). The applied power () was determined directly from Joule heating in the tip under the applied current using Eq.(11). The resulting data were fitted using linear regression, yielding an value of K/W as shown in Fig. 3. An instrumental uncertainty of 0.1 K in due to laser illumination was considered in the analysisspiece2019quantitative. It should be noted that the uncertainty associated with laser illumination may be relatively large; however, since the analysis focuses on the relative change in temperature (), the impact of this uncertainty is reduced. This result is consistent with previously reported values for similar probes, thereby validating the accuracy of the proposed measurement technique huang2024violation; razeghi2023single.
Here, denotes the voltage drop across the tip, extracted from the known SThM probe resistance and the Wheatstone bridge input voltage (Fig. 1(a)).